# Catalog of Mock Stellar Streams in Milky Way-like Galaxies

### v1

@@ 3.3 Stream Photometric Predictions
We integrate our PDFs to get cumulative distributions of the number of streams that contain $N_{\rm stars}$ observable stars for a given instrument, for a given halo: $N(>N_{\rm stars})$. This result is shown in Figure 7. For each instrument, the region plotted encapsulates all four individual distributions for each halo. The uncertainty arises from the differing number of total streams in each halo. We also plot the theoretical limiting distribution if every stream star in the catalog were observable. The figure shows how impactful LSST and Roman will be in identifying new streams. For example, Gaia is only able to observe at least $10^{4}$ stars in about $60$ streams, but LSST and Roman would be able to observe the same number of stars in hundreds of streams. Roman would be able to detect at least 1000 stars in almost the entire sample across all four halos.
In Figure 8 we divide the streams into three approximately equal sized-groups based on their present-day galactocentric radius, for the adopted LSST magnitude limit of $g<24.5$. Even for streams beyond 12 kpc, LSST will have resolving power to identify several dozen, up to potentially more than one hundred member stars according to our catalog, depending on the true intrinsic radial distribution of streams in the outer halo. An extended catalog of identified GC stellar streams at large galactocentric radii will be crucial for better understanding the difference between baryonic and DM perturbations to streams, which will be key if streams are to one day probe the subhalo mass function and constrain the nature of DM.
Figure 8: Number of observable stars per stream binned by their present-day galactocentric radii. Each shaded region encapsulates the range between all four halos. For example, the number of streams beyond 12 kpc with 100 or more observable stars ranges from 6018060-180 depending on the halo. The uncertainty arises primarily from the varying number of streams in each halo. Depending on the true number, our catalogs predict there are anywhere from $\sim$60 to well over one hundred streams beyond 12 kpc with 1000 or more stars below the detection limit of LSST.
@@
Table 1 lists the lowest mass of stream stars detectable by the three instruments in each bin of galactocentric radius from Figure 8. For each stream in each halo, we compute the lowest mass observable according to each instrument’s limiting magnitude, then for each bin we compute the 25th and 75th percentile values. The table displays these values averaged across the four halos. Stars in the middle bin are typically at the closest heliocentric distances, which will allow LSST and Roman to sample the mass function even below $0.2\,{\mathrm{M}_{\odot}}$. Gaia is also relatively effective in this radius bin, being able to probe down to $\sim 0.3\,{\mathrm{M}_{\odot}}$ for a quarter of the streams within this galactocentric distance range. However, in the outer bin, the Gaia limit increases dramatically. The percentiles reported here are only for the streams in which there is more than one star below the detection limit of Gaia; in every halo, there are streams that Gaia misses entirely. The exact number varies according to the varying quantity and radial distribution of streams in each halo, from a minimum of 4 in 523889, to a maximum of 91 in 519311. The FIRE halo m12w, which most closely matches the MW in terms of its GC radial distribution (see Appendix C), there are 19 streams for which Gaia can observe no stars. Given the fact that at minimum several dozen stars are needed to identify a stream with current methods (e.g., Malhan & Ibata, 2018), this result suggests that there are anywhere from several dozen up to over one hundred streams in the MW halo undetectable with Gaia.
@@ 4 Discussion
Our model catalog of 1473 tidal features from disrupted GCs across four cosmological halos contains a diverse set of objects including hundreds of thin streams, which can be used for many purposes such as analysis of a stream population in a MW-like galaxy, mock photometry of individual streams, modeling of streams to constrain the Galactic potential, and testing stream detection algorithms. Here we describe additional possible uses of the catalog, analyze our initial results further, and discuss possible limitations.
@@ 4.1 Orbital Properties of Thin and Wide Streams
Given the orbital information of the GC progenitors for the last several Gyr, we can further distinguish between the types of orbits that form long, thin stellar streams and those that do not. In Figure 9 we show the energy-angular momentum diagram for all 287 progenitor orbits for the features in halo 523889. In our coordinate system prograde orbits are on the right (positive $L_{z}$) and retrograde orbits are on the left (negative $L_{z}$). The majority of thin streams based on our definition have energies greater than $0.1\penalty 10000\ \rm kpc^{2}\penalty 10000\ Myr^{-2}$, indicating that orbits more strongly bound than this (thereby residing closer to the Galactic center) are heated to the point where they do not form thin coherent streams. However, there are also many streams on prograde circular orbits which remain thin deeper into the potential well. Notably, the current known phase-space distribution of GC streams in the MW (Ibata et al., 2024) does not have such streams. It is possible that these streams have evaded detection thus far; they are closer to the Galactic center and appear on the sky in regions of higher background stellar density, which makes them harder to detect. Further analysis of the orbits in connection with the resulting tidal features should help to clarify what types of streams to expect on what types of orbits.
@@ 4.2 Impact of Future Observatories
Our findings agree with previous results that both LSST and Roman will greatly expand our knowledge of streams in the MW and other galaxies in the Local Group (Pearson et al., 2019; Shipp et al., 2023; Aganze et al., 2024; Pearson et al., 2024). Specifically, we find that the number of observable stars per stream will increase drastically.
We can also use these data to estimate the number of streams potentially observable at present that have evaded detection. Using our adopted definition of a thin stream, there are between 60 and 124 thin streams with 100 or more observable stars with Gaia in each halo, compared with the 80 or so known GC stellar streams. This suggests the possibility of several dozen missing streams within the detection limit of Gaia. However, the number may even be greater than this, since we are only tracking streams which formed in the last $3.5$ Gyr (see Pearson et al. 2024). One possible reason for evading detection is extinction from MW dust near the Galactic plane. Out of our set of thin streams with at least 100 stars below the Gaia detection limit, the total number with progenitor locations on the sky with $|b|<10^{\circ}$ ranges between 36 – 64 for our four different halos. These streams would be partially or completely obscured in optical wavelengths due to foreground dust. A future infrared survey astrometry telescope could make it possible to detect such streams, such as the proposed GaiaNIR mission (Hobbs & Høg, 2018).
@@ 4.3 Synthetic Photometry of Individual Streams
It is also possible to analyze individual streams in the catalog and create analogs to real observed systems. An example is in Figure 10, which shows the color-magnitude diagram (CMD) of a stream with similar mass and orbit to Pal 5. Its progenitor has an orbital pericenter of $11$ kpc, apocenter of $21$ kpc, and a present-day mass of $1.6\times 10^{4}\,{\mathrm{M}_{\odot}}$. In order to match the observed properties of Pal 5, for this Figure only we keep the Sun fixed at $8.122$ kpc from the Galactic center, but perform a rotation in the xyx-y plane to best match Pal 5’s radial velocity ($58.6\penalty 10000\ \rm km\penalty 10000\ s^{-1}$; Baumgardt et al., 2019) and heliocentric distance ($21.9$ kpc; Baumgardt & Vasiliev, 2021). The optimal rotation gives our mock cluster a heliocentric distance of $20.9$ kpc and $r_{v}=-64.8\penalty 10000\ \rm km\penalty 10000\ s^{-1}$. We use the MIST isochrone for an age of $10.355$ Gyr and metallicity of $\mathrm{[Fe/H]}=-1.56$ to compute mock SDSS $g$ and $r$ magnitudes for the stars with their updated distances, assuming no extinction along the line of sight.
Stars along the stream have slightly different heliocentric distances than the progenitor, which creates measurable scatter around the original isochrone. We treat the mock CMD as if it were observed data and compute the best-fitting isochrone for all stars with $g<24.5$ to the widened CMD. We sample a $100\times 100$ grid of ages and metallicities ranging from $\{9,13\}\penalty 10000\ \rm Gyr$ and $\{-2,-0.5\}\penalty 10000\ \rm dex$, and minimize the summed squared distance from each point to the isochrone. The best fitting isochrone from this search has an age of $9.2\penalty 10000\ \rm Gyr$ and $\rm Fe/H]=-1.23$. The two curves are similar near the main sequence turnoff but begin to deviate by about 0.1 dex at $r\approx 25$, where the stars were not included in the fitting. There is also a notable discrepancy on the red giant branch where there are fewer stars. Notably, if all stars are included, we get a much closer fit giving an age of $9.97$ Gyr and $\rm[Fe/H]=-1.49$. This example shows that deeper imaging, such as that of Roman, will allow for significantly better constraints on a stream progenitor’s age and metallicity given its CMD.
@@
len
Figure 10: Color-magnitude diagram for a stream with similar apocenter and pericenter values to Pal 5, color-coded by stellar mass. The corresponding cluster in the catalog has an age of $10.36$ Gyr and we set the metallicity to $\mathrm{[Fe/H]}=-1.56$. The MIST isochrone for a stellar population with this age and metallicity at a distance of $20.9$ kpc is shown as a dashed line. Next to it is a solid line, representing the best-fit isochrone from the data with $g<24.5$ ($9.2$ Gyr, $\mathrm{[Fe/H]}=-1.23$). The scatter of stars is due to stream stars having different distances from Earth. The dotted lines located at $\pm 0.1$ dex in color from the best-fit isochrone show the bounds needed to capture 80% of all stream members.
@@ 4.4 Connection to Host Potential
Another promise of streams is studying the connection between the stream population and the assembly history of the host halo (Carlberg, 2017; Bonaca et al., 2021; Li et al., 2022). Our GC formation and stream models are based on cosmologically-formed halos, for which we know the true assembly history. Such a system has proved valuable for reconstruction of the galactic assembly histories with GCs (Chen & Gnedin, 2024b). With our stream catalog, a more comprehensive analysis can be applied to better understand the MW assembly history.
@@ 4.6 Modeling Limitations
Selecting galaxies from cosmological simulations for our model calculation leads to two numerical limitations. First, we are limited by both the spatial and temporal resolution of the simulations. Since our model GC initial positions are assigned to particles in TNG and FIRE, they can be limited by each simulation’s force-softening length. In FIRE, this is only $4$ pc for stellar particles, but for TNG it is $\sim 300$ pc, which can lead to errors in the distribution of GCs in the inner galaxy at a lookback time of $3.5$ Gyr. Moreover, the time resolution between snapshots used in the potential fitting varies between 100 and 200 Myr for TNG and 10 Myr - 1 Gyr for FIRE, which washes out important baryonic components of the potential that can impact streams, such as the spiral arms and the Galactic bar. There is also no LMC analog in our selected halos, which has been shown to leave significant perturbations on streams in its vicinity (e.g., Brooks et al., 2025). Higher resolution large-scale simulations will be needed to model all of these affects more accurately and study their impact on the total Galactic stream population.
Second, the fitting of the potential using BFE has its own limitations when considering the appropriate choices of $l_{\rm max}$ and $m_{\rm max}$. Selecting too low a value for these parameters, particularly $l_{\rm max}$, will smooth out too much of the halo’s features and substructure. However, too high a value of $l_{\rm max}$ will cause the potential to be overfit in regions where the particle numbers are low. This is especially true close to the Galactic center, where the number of particles enclosed within a spherical shell becomes small. For example, out of all the particles present at the last snapshot for TNG halo 523889, only $3\%$ are within 1 kpc of the Galactic center, and only $11\%$ are within 4 kpc, which roughly corresponds to the scale length of the MW disk. At the same time, it is important to model the potential accurately at small galactocentric radii, since the cluster mass-loss on eccentric orbits is dominated at pericenter. We made the selection of $l_{\rm max}$ and $m_{\rm max}$ after careful consideration of both competing issues (see Appendix A), but further improvements in this area will be essential for high-precision modeling of streams in time-evolving potentials.

### v2

@@ 3.3 Stream Photometric Predictions
We integrate our PDFs to get cumulative distributions of the number of streams that contain $N_{\rm stars}$ observable stars for a given instrument, for a given halo: $N(>N_{\rm stars})$. This result is shown in Figure 7. For each instrument, the region plotted encapsulates all four individual distributions for each halo. The uncertainty arises from the differing number of total streams in each halo. For Roman and LSST, we account for sky coverage by scaling each CDF according to what percent of the sky the instrument is expected to cover: $65\%$ for LSST, and $12\%$ for Roman’s High-Latitude Wide-Area Survey (although in the future Roman is expected to observe other portions of the sky as well). We also plot the theoretical limiting distribution if every stream star in the catalog were observable. The figure shows how impactful LSST and Roman will be in identifying new streams. For example, Gaia is only able to observe at least $10^{4}$ stars in about $60$ streams, but LSST would be able to observe the same number of stars in hundreds of streams. Roman would be able to detect at least 1000 stars in 304030-40 streams, which corresponds to essentially every stream in its available region of sky across all four halos.
In Figure 8 we divide the streams into three approximately equal sized-groups based on their present-day galactocentric radius, for the adopted LSST magnitude limit of $g<25$. We once again scale each CDF by $65\%$ to account for LSST’s sky coverage in the Southern Hemisphere. Even for streams beyond 12 kpc, LSST will have resolving power to identify several dozen, up to potentially more than one hundred member stars per stream according to our catalog, depending on the true intrinsic radial distribution of streams in the outer halo. An extended catalog of identified GC stellar streams at large galactocentric radii will be crucial for better understanding the difference between baryonic and DM perturbations to streams, which will be key if streams are to one day probe the subhalo mass function and constrain the nature of DM.
Figure 8: Number of streams with $N_{\rm stars}$ or more observable stars by LSST binned by their present-day galactocentric radii. Each shaded region encapsulates the range between all four halos. For example, the number of streams beyond 12 kpc with 100 or more observable stars ranges from 40 to over 100 depending on the halo. The uncertainty arises primarily from the varying number of streams in each halo. This predicts that LSST has the potential to greatly expand our census of known streams beyond the stellar disk.
@@
Table 3 lists the lowest mass of stream stars detectable by the three instruments in each bin of galactocentric radius from Figure 8. For each stream in each halo, we compute the lowest mass observable according to each instrument’s limiting magnitude, then for each bin we compute the 25th and 75th percentile values. The Table displays these values averaged across the four halos. Stars in the middle bin are typically at the closest heliocentric distances, which will allow LSST and Roman to sample the mass function even below $0.2\,{\mathrm{M}_{\odot}}$. Gaia is also relatively effective in this radius bin, being able to probe down to $\sim 0.3\,{\mathrm{M}_{\odot}}$ for a quarter of the streams within this galactocentric distance range. However, in the outer bin, the Gaia limit increases dramatically. The percentiles reported here are only for the streams in which there is more than one star below the detection limit of Gaia; in every halo, there are streams that Gaia misses entirely. The exact number varies according to the varying quantity and radial distribution of streams in each halo, from a minimum of 4 in 523889, to a maximum of 91 in 519311. The FIRE halo m12w, which most closely matches the MW in terms of its GC radial distribution (see Appendix A), has 19 streams for which Gaia can observe no stars. Given the fact that at minimum several dozen stars are needed to identify a stream with current methods (e.g., Malhan and Ibata, 2018), this result suggests that there are anywhere from several dozen up to over one hundred streams in the MW halo undetectable with Gaia.
@@ 4 Discussion
Our model catalog of 1471 tidal features from disrupted GCs across four cosmological halos contains a diverse set of objects including hundreds of thin streams, which can be used for many purposes such as analysis of a stream population in a MW-like galaxy, mock photometry of individual streams, modeling of streams to constrain the Galactic potential, and testing stream detection algorithms. Here we describe additional possible uses of the catalog, analyze our initial results further, and discuss possible limitations.
@@ 4.1 Orbital Properties of Thin and Wide Streams
Given the orbital information of the GC progenitors for the last several Gyr, we can further distinguish between the types of orbits that form long, thin stellar streams and those that do not. In Figure 9 we show the energy-angular momentum diagram for all 287 progenitor orbits for the features in halo 523889. In our coordinate system prograde orbits are on the right (positive $L_{z}$) and retrograde orbits are on the left (negative $L_{z}$). Over 90 percent of thin streams based on our definition have energies greater than $0.13\penalty 10000\ \rm kpc^{2}\penalty 10000\ Myr^{-2}$, indicating that orbits more strongly bound than this (thereby residing closer to the Galactic center) are heated to the point where they do not form thin coherent streams. This is also influenced by the fact that these strongly bound orbits have much smaller orbital periods, meaning they have completed more orbits within our integration time than the other clusters. While most of these strongly bound orbits form wide streams, there are also many streams on prograde circular orbits which remain thin deeper into the potential well. Notably, the current known phase-space distribution of GC streams in the MW (Ibata et al., 2024) does not have such streams. It is possible that these streams have evaded detection thus far; they are closer to the Galactic center and appear on the sky in regions of higher background stellar density, which makes them harder to detect. It is also possible that true streams on these orbits are susceptible to heating from disk properties which are not modeled in this work. Further analysis of the orbits in connection with the resulting tidal features should help to clarify what types of streams to expect on what types of orbits.
@@ 4.2 Impact of Future Observatories
Our findings agree with previous results that both LSST and Roman will greatly expand our knowledge of streams in the MW and other galaxies in the Local Group (Pearson et al., 2019; Shipp et al., 2023; Aganze et al., 2024; Pearson et al., 2024). Specifically, we find that the number of observable stars per stream will increase drastically. This provides the chance to both significantly increase the amount of confirmed members of known streams, and shows potential to find entirely new streams previously undetectable. While we have not modeled a MW background population in this work, which is a major hurdle in the detection of new streams, Pearson et al. (2024) argues that streams whose main sequence turnoff is one magnitude below the detection limit should be observable based on previous stream discoveries (Shipp et al., 2018). Applying this to our dataset, we find that every stream in the entire catalog could in principle be detected by Roman. For LSST, we find that after accounting for its footprint on the sky that an average of 63% of streams are detectable across the hour halos. Given that LSST can only image 65% of the sky, this means that given this definition of detectability, it will be able to detect nearly every stream in the Southern Hemisphere.
We can also use these data to estimate the number of streams potentially observable at present that have evaded detection. Using our adopted definition of a thin stream, there are between 60 and 124 thin streams with 100 or more observable stars with Gaia in each halo, compared with the 80 or so known GC stellar streams. This suggests the possibility of several dozen missing streams within the detection limit of Gaia. However, the number may even be greater than this, since we are only tracking streams which formed in the last $3.5$ Gyr (see Pearson et al. 2024). One possible reason for evading detection is extinction from MW dust near the Galactic plane. Out of our set of thin streams with at least 100 stars below the Gaia detection limit, the total number with progenitor locations on the sky with $|b|<10^{\circ}$ ranges between 36 – 64 for our four different halos. These streams would be partially or completely obscured in optical wavelengths due to foreground dust. A future infrared survey astrometry telescope could make it possible to detect such streams, such as the proposed GaiaNIR mission (Hobbs and Høg, 2018).
We acknowledge that a true detailed analysis on the observability of streams for a given instrument would require modeling the background MW population and extinction from Galactic dust. There is also the matter of what data is available per instrument: for example, Gaia also computes proper motions of potential members, which can greatly enhance the ability to detect streams even with its lower magnitude limit. However, LSST will also produce proper motions for stars over a 10-year baseline, and obtaining proper motions with Roman is also possible in the future. Considering these facts, we predict from our findings that LSST and Roman will expand the known number of stellar streams and give us better insight into previously known streams.
@@ 4.3 Synthetic Photometry of Individual Streams
It is also possible to analyze individual streams in the catalog and create analogs to real observed systems. An example is in Figure 10, which shows the color-magnitude diagram (CMD) of a stream with similar mass and orbit to Pal 5. Its progenitor has an orbital pericenter of $11$ kpc, apocenter of $21$ kpc, and a present-day mass of $1.6\times 10^{4}\,{\mathrm{M}_{\odot}}$. In order to match the observed properties of Pal 5, for this Figure only we keep the Sun fixed at $8.122$ kpc from the Galactic center, but perform a rotation in the xyx-y plane to best match Pal 5’s radial velocity ($58.6\penalty 10000\ \rm km\penalty 10000\ s^{-1}$; Baumgardt et al., 2019) and heliocentric distance ($21.9$ kpc; Baumgardt and Vasiliev, 2021). The optimal rotation gives our mock cluster a heliocentric distance of $20.9$ kpc and $v_{r}=-64.8\penalty 10000\ \rm km\penalty 10000\ s^{-1}$. We use the MIST isochrone for an age of $10.355$ Gyr and metallicity of $\mathrm{[Fe/H]}=-1.56$ to compute mock SDSS $g$ and $r$ magnitudes for the stars with their updated distances, assuming no extinction along the line of sight.
Stars along the stream have slightly different heliocentric distances than the progenitor, which creates measurable scatter around the original isochrone. We treat the mock CMD as if it were observed data and compute the best-fitting isochrone for all stars with $g<25$ to the widened CMD. We sample a $100\times 100$ grid of ages and metallicities ranging from $\{9,13\}\penalty 10000\ \rm Gyr$ and $\{-2,-0.5\}\penalty 10000\ \rm dex$, and minimize the summed squared distance from each point to the isochrone. The best fitting isochrone from this search has an age of $9.2\penalty 10000\ \rm Gyr$ and $\rm Fe/H]=-1.23$. The two curves are similar near the main sequence turnoff but begin to deviate by about 0.1 dex at $r\approx 25$, where the stars were not included in the fitting. There is also a notable discrepancy on the red giant branch where there are fewer stars. Notably, if all stars are included, we get a much closer fit giving an age of $9.97$ Gyr and $\rm[Fe/H]=-1.49$. This example shows that deeper imaging, such as that of Roman, will allow for significantly better constraints on a stream progenitor’s age and metallicity given its CMD.
@@
Figure 10: Color-magnitude diagram for a stream with similar apocenter and pericenter values to Pal 5, color-coded by stellar mass. The corresponding cluster in the catalog has an age of $10.36$ Gyr and we set the metallicity to $\mathrm{[Fe/H]}=-1.56$. The MIST isochrone for a stellar population with this age and metallicity at a distance of $20.9$ kpc is shown as a dashed line. Next to it is a solid line, representing the best-fit isochrone from the data with $g<25$ ($9.2$ Gyr, $\mathrm{[Fe/H]}=-1.23$). The scatter of stars is due to stream stars having different distances from Earth. The dotted lines located at $\pm 0.1$ dex in color from the best-fit isochrone show the bounds needed to capture 80% of all stream members.
@@ 4.4 Connection to Host Potential
Another promise of streams is studying the connection between the stream population and the assembly history of the host halo (Carlberg, 2017; Bonaca et al., 2021; Li et al., 2022). Our GC formation and stream models are based on cosmologically-formed halos, for which we know the true assembly history. Such a system has proved valuable for reconstruction of the galactic assembly histories with GCs (Chen and Gnedin, 2024b). With our stream catalog, a more comprehensive analysis can be applied to better understand the MW assembly history.
@@ 4.6 Modeling Limitations
Selecting galaxies from cosmological simulations for our model calculation leads to two numerical limitations. First, we are limited by both the spatial and temporal resolution of the simulations. Since our model GC initial positions are assigned to particles in TNG and FIRE, they can be limited by each simulation’s force-softening length. In FIRE, this is only $4$ pc for stellar particles, but for TNG it is $\sim 300$ pc, which can lead to errors in the distribution of GCs in the inner galaxy at a lookback time of $3.5$ Gyr. Moreover, the time resolution between snapshots used in the potential fitting varies between 100 and 200 Myr for TNG and 10 Myr - 1 Gyr for FIRE, which washes out important baryonic components of the potential that can impact streams, such as the spiral arms and the Galactic bar. There is also no LMC analog in our selected halos, which has been shown to leave significant perturbations on streams in its vicinity (e.g., Koposov et al., 2023; Lilleengen et al., 2023; Brooks et al., 2025). Higher resolution large-scale simulations will be needed to model all of these affects more accurately and study their impact on the total Galactic stream population.
Second, the fitting of the potential using our BFE techniques in agama has its own limitations when considering the appropriate choices of $l_{\rm max}$ and $m_{\rm max}$. Selecting too low a value for these parameters, particularly $l_{\rm max}$, will smooth out too much of the halo’s features and substructure. However, too high a value of $l_{\rm max}$ will cause the potential to be overfit in regions where the particle numbers are low. This is especially true close to the Galactic center, where the number of particles enclosed within a spherical shell becomes small. For example, out of all the particles present at the last snapshot for TNG halo 523889, only $3\%$ are within 1 kpc of the Galactic center, and only $11\%$ are within 4 kpc, which roughly corresponds to the scale length of the MW disk. At the same time, it is important to model the potential accurately at small galactocentric radii, since the cluster mass loss on eccentric orbits is dominated at pericenter. We made the selection of $l_{\rm max}$ and $m_{\rm max}$ after careful consideration of both competing issues (see Appendix B), but further improvements in this area will be essential for high-precision modeling of streams in time-evolving potentials.
In addition to these limitations, our prescription for the mass loss of GCs is further limited by the fact that Equation 2 has no dependence on the internal dynamics of the cluster itself. While our model is calibrated from the N-body models of Gieles and Gnedin (2023) for GCs with typical density profiles, clusters which are more extended or more densely concentrated will have different mass-loss rates. Furthermore, recent work has shown that a population of stellar mass black holes within the GC can affect both the mass-loss rate of GCs as well as the properties of stellar streams (Gieles and Gnedin, 2023; Roberts et al., 2025; Weatherford and Bonaca, 2025). A mass-loss rate calculation which considers internal cluster dynamics as well as a higher resolution in the snapshot cadence of the underlying potential can more accurately capture the physics which causes gaps and density variations along stream tracks. This type of modeling will be important for understanding the differences between baryonic and dark matter perturbations along stream tracks.